OpenStudy (anonymous):

12 of the 40 cars pulled by by the locomotive were tankers. A: What fraction of the cars were tankers B: What percent of the cars were not tankers

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

Very simple. Total vehicles = 40. Fraction of cars = cars / total = 12 / 40 = 30 percent or 3/10ths. Then, the percentage of vehicles that were not tankers is 100% - 30% or 1 - 0.3 = 70% or 0.7.

OpenStudy (anonymous):

ok thanks i understand now